An airplane is flyigg at 195m/s due North makes a gradual turn following a circular path to fly due south. the turn takes 45 seconds to complete
a) what is teh radius of the curve that the plane follows in making the turn (2.79E3M answer)
b) what is teh magnitude of the centripetal acceleration during the turn (13.63m/s answer)
pls can u show me the approach, im stuck
i used the equations v=2*pi*r/t and set it equal to r??

Answer by htmentor(1343)

v = w*r where w is the angular velocity and r is the radius
v = 195 m/s
The plane completes a half circle in 45 s.
Since a half circle is equal to pi radians, we have:
r = 195/(pi/45) = 2793 m
The centripetal acceleration is v^2/r = 195^2/2793 = 13.63 m/s^2